Teachers divided students into groups of three each group of three wrote a report that had nine pictures in it the students Used 558 pictures altogether how many students were there all together
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

Teachers divided students into groups of three each group of three wrote a report that had nine pictures in it the students Used 558 pictures altogether how many students were there all together.
---------
# of students = x
# of groups = x/3
---
# of pictures = (x/3)9 = 558
3x = 558
x = 186
--------------
Cheers,
Stan H.
-------------

Answer by richwmiller(17219)

The teachers divided the students into groups of three. Each group of three wrote a report that had nine pictures in it. The students used 558 pictures altogether. How many students were there all together ?
558/9=62 groups of three
62*3=186
or
3*558/9=558/3=186
There were 186 students. There were 558 pictures used by 62 groups of three